if 2 raise to the power x = 12 raise to the power x =3 raise to the power y
then, x+2y/xy=?
1
Expert's answer
2013-02-01T09:46:54-0500

if 2 raise to the power x = 12 then x = 3.585
if 2 raise to the power y =3 then y = 1.585
then x+2y/xy = 3.585 +2*1.585/1.585*3.585 = 4.14288